OpenStudy (anonymous):

Use elimination to solve the system and . What is the x-value in the solution?

mathslover (mathslover):

may i have the given expression please

OpenStudy (anonymous):

sorry! 3x-5y=8 and 4x-7y=10

mathslover (mathslover):

no problem : 3x-5y = 8 .. can u solve for x here ?

mathslover (mathslover):

3x = 8 + 5 y now @jasminecrump18 ?

OpenStudy (anonymous):

so subtract 8 from both sides?

OpenStudy (anonymous):

you have to subtract both equations

OpenStudy (anonymous):

Multiply first equation 4 with and second equation with 3..

OpenStudy (anonymous):

multiply eqn i by 4 and eqn ii by 3 and subtract eqn ii from eqn i. you will get the value of y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Or if you want to find x directly then you must multiply 7 to first equation and 5 to second equation...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Show your work after multiplication..

OpenStudy (anonymous):

ok thanks